OpenStudy (anonymous):

Rhonda deposited $4,227.29 into a savings account with an interest rate of 4.9% compounded monthly. About how long will it take for the account to be worth $9,000?

OpenStudy (anonymous):

please help me you guys

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I can help you

OpenStudy (anonymous):

33 years, 2 months 3 years, 8 months 15 years, 10 months 15 years, 6 months

OpenStudy (anonymous):

thts the choices

OpenStudy (tkhunny):

Have you considered the standard accumulation formula? \(4,227.29\left(1+\dfrac{0.049}{12}\right)^{12*n} = 9000\) Solve for n, the number of years. However, as it asks for ABOUT how long, I'm tempted to use the "Rule of 72"! 9000/4227.29 = 2.129 -- That's pretty close to 2. 72/4.9 = 14.7 is a little short. So 15 years? This doesn't do a lot for us,. but it does rule out the first two choices.

OpenStudy (jim766):

15 yrs 10 mths

OpenStudy (anonymous):

thk you <3